OpenStudy (anonymous):

Suppose the point (2, -7) is on the graph of y = f(x). Find the corresponding point on the following transformations of f. y = −13f(x + 3)

OpenStudy (helder_edwin):

u know that f(2)=-7

OpenStudy (anonymous):

so what are the coordinates of y=-13f(x+3)

OpenStudy (helder_edwin):

first x+3=2 then x=-1

OpenStudy (helder_edwin):

so if x=-1 then \[ \large y=-13\cdot f(-1+3)=-13\cdot f(2)=-13\cdot(-7)=91 \]

OpenStudy (anonymous):

so (-1, 91)

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

oh thank you so much

OpenStudy (helder_edwin):

y r welcome